# Enhancing High-Level Synthesis with Automated Pragma Insertion and Code Transformation Framework

### v3

@@ 4.3. Objective Function
We structure the objective function akin to the approach described in (Pouget et al., 2024b, a), tailoring it to meet the specific needs of our problem. Here, $Lat_{2}$ corresponds to the fine-grained unrolled tile, $Lat_{1}$ denotes the pipeline tile that incorporates the fine-grained unrolled tile, and recursively, $Lat_{0}$ encompasses $Lat_{1}$, enabling coarse-grained unrolling and facilitating on-chip data caching.
@@
$\left.\begin{cases}&Lat_{2}=IL_{par}+IL_{seq}\times\prod_{l\in\mathcal{L}^{red}}\times\log_{2}(TC_{l,2})\\ &Lat_{1}=Lat_{2}+II\times(TC_{l,1}-1)\\ &Lat_{0}=\prod_{l\in\mathcal{L}}\frac{TC_{l,0}}{loop_{l}\_UF}\times Lat_{1}\\ &L_{mem}=\sum_{l\in\mathcal{L}}\max_{a\in\mathcal{A}}(cache_{l}\_arr_{a}\times ft\_array_{a}\_loop_{l})\\ &obj\_func=Lat_{0}+L_{mem}\end{cases}\right.$
5. Code Generation and Optimization
The NLP allows us to find the trip count of each loop, the array partitioning for each array, and the size of the on-chip buffers. By considering the loop order and the pragmas applied to each loop as described in Section 3, we can directly generate the code. We automatically generate the functions that load data from off-chip to on-chip with the maximum burst size possible and vice versa with the functions that store the data off-chip. Additionally, we incorporate further optimizations to enhance the QoR.
5.1. Optimization for non-constant trip count
To optimize loops with non-constant trip counts, we implement a code transformation that preserves the NLP’s estimation while simplifying compilation for the HLS compiler. We achieve this by replacing loops with non-constant trip counts with the maximal trip count computed using PoCC (PoCC, [n. d.]). Subsequently, we create a function for all statements iterated by this loop, ensuring compliance with the constraints of the non-constant trip count loop. This function returns the computation if the constraints are met or returns just the value of the output otherwise. Introducing a condition solely above the statement results in excessive compilation times. Therefore, we employ these techniques to reduce compilation time and achieve designs with a good QoR.
5.2. Overlapping computation and communication
The constraints of the NLP consider the sum of the footprints of all arrays used during the design execution. Thus, we have the option, while already anticipating resource constraints, to overlap the transfer of arrays between off-chip and on-chip (or vice versa) at the same level or between the transfer of the arrays needed for the next loop body with the computation of the current loop body.
To achieve this, we separate each loop body into different functions, as well as the functions that transfer input data between off-chip and on-chip. Additionally, we create new functions that execute these independent functions in parallel. To ensure parallel execution, within these newly created functions designed for overlap, we include only independent functions.
5.3. Code transformation for HLS
To simplify the understanding of the HLS compiler, particularly to ensure that the compiler can pipeline loops with the correct initiation interval (II), we simplify the reductions (regardless of whether we use tree reductions) in the fine-grained unrolled part. As specified in section 3, the fully unrolled loops are placed innermost. Therefore, before the reduction, we create a variable to accumulate the reduction, which we then add to the output.
Additionally, we incorporate the pragma loop_flatten off on all reduction loops above the pipeline. While theoretically, this adjustment could enhance the QoR, in practice, the compiler might choose not to pipeline at all if it cannot flatten the loop and pipeline correctly.
@@ 6.1. Setup
We employ kernels sourced from Polybench/C 4.2.1 (PolyBench, [n. d.]), supplemented by a Convolutional Neural Network (CNN) kernel and matrix multiplication tasks akin to those in the BERT transformer model. Our computations utilize single-precision floating-point data types as the default, facilitating comparison with AutoDSE and NLP-DSE. We operate on medium-sized datasets from PolyBench/C (PolyBench, [n. d.]). We also compare our approach with HARP (Sohrabizadeh et al., 2023), which seeks to enhance High-Level Synthesis design exploration by integrating a Graph Neural Network model that predicts HLS tool behavior. They conduct exploration within this space utilizing the GNN model for one hour and then synthesize the top-10 designs. We use the medium-sized Polybench kernels found in their training set with double-precision floating-point data types, aiming to deploy HARP under optimal conditions. As detailed in (Pouget et al., 2024b, a), HARP’s effectiveness is not universally applicable without the necessary fine-tuning or training, particularly when confronted with diverse problem sizes.
We chose these problem sizes to demonstrate the efficacy of our approach in transforming codes and inserting pragmas in scenarios where fully unrolling is not feasible. Additionally, we aim to illustrate the effectiveness of our framework in tiling code for substantial tasks such as CNN and bert_3072_100. This approach ensures that only a portion (a tile) of each array fits on-chip, aligning with the available memory constraints. The problem size and original loop order of CNN are I,J=256, H,W=224 and P,Q=5. For bert_n_m matrix multiplication the schedule and problem size are I=n, J=m, K=n (reduction).
We utilize the AMD/Xilinx Vitis HLS compiler (Xilinx, 0232) to showcase the effectiveness of Sisyphus. Report generation is performed using AMD/Xilinx Vitis 2023.2. When enabling tree reduction, we utilize the ”funsafe math optimizations” option to enable commutative/associative reduction operators, thereby facilitating reductions to be implemented in logarithmic time. In our comparison with HARP (Sohrabizadeh et al., 2023), we utilized the same parameters as the authors, employing Vitis 2021.1 and activating the option for tree reduction. As our designated hardware platform, we deploy the Xilinx Alveo U200 device with a target frequency of 250 MHz. We conduct an analysis of the kernels and automatically generate each NLP problem using PoCC (PoCC, [n. d.]). Solving these NLP problems, we employ the AMPL description language and utilize the commercial BARON solver version 21.1.13 (Tawarmalani and Sahinidis, 2005; Sahinidis, 2017). For our experimental setup, we harness the computing power of 2 AMD EPYC 7V13 64-Core Processor.
@@ 6.2. Experimental Evaluation
The aim of this evaluation is to demonstrate our capability to perform the proper transformation while simultaneously inserting the correct hardware directives. To achieve this, we compare our approach with three frameworks, that conduct design space exploration to insert pragmas with a fixed schedule.
The input codes provided are the original codes from Polybench or as described in the subsection 6.1. However, we modify the kernels containing loops with non-constant trip counts using the method outlined in Section 5.3. This adjustment ensures a fair evaluation with the transformations we apply.
We compare our method with AutoDSE (Sohrabizadeh et al., 2021) and NLP-DSE (Pouget et al., 2024a, b) described in Section 2, and HARP (Sohrabizadeh et al., 2023).
We generate the AutoDSE space automatically by running the command ds_generator. We replace the unroll factor and tile size with all factors of the trip count to ensure consistency within the space. However, since maximum array partitioning is not a parameter within AutoDSE’s space, this space, as observed later with the kernel atax, is slightly larger. AutoDSE is executed using the bottleneck method with a DSE timeout set to 1,000 minutes and a timeout of 180 minutes for each HLS synthesis. NLP-DSE is executed with the parameters specified in the paper, including a 30-minute timeout for the BARON solver and the space provided as input is $\{\infty,2048,1024,512,256,128,64,32,16,8,1\}$. For HARP, we adhere to the parameters outlined by the authors mentioned earlier.
In evaluating Sisyphus, we set a solver timeout of 14,400 seconds (4 hours) for the Baron solver, which generates only one design. By default, we assume that the compiler can efficiently reuse resources between statements that are not executed in parallel. If the synthesis results in the design exceeding DSP utilization, we relaunch the NLP with a pessimistic DSP utilization assumption—meaning we consider no reuse between statements—and regenerate the design accordingly. For heat-3d, symm, syrk, syr2k, and bert_3072_100 we had to apply the constraint indicating no reuse (Eq. 16), as the constraint with optimistic reuse (Eq. 15) resulted in kernels with resource over-utilization. The maximum partitioning constraint, $max_{part}$ (cf. Section 4), is set at 1024, as determined by our use of AMD/Xilinx compilers and FPGAs, adhering to their specified limit.
Each synthesis process begins with a C-simulation to verify the correctness of the generated code.
@@ 6.3. Comparison of the methods
6.3.1. AutoDSE and NLP-DSE
Table 3 provides a comparison among AutoDSE, NLP-DSE, and Sisyphus. The TR column denotes the status of the tree reduction option, while the T column indicates the specific transformations applied, with D representing distribution, T for data tiling, and P for permutation. Across all evaluated kernels, we consistently achieve comparable or superior performance, with the exception of the atax kernel, where AutoDSE yields designs approximately 1% to 2% faster, and heat-3d, where NLP-DSE makes a 1% improvement. However, the designs obtained by AutoDSE fall outside our defined space due to the stringent constraint of a maximum array partitioning limit of 1024 imposed by AMD/Xilinx Vitis. The design found by NLP-DSE for heat-3d shares the same schedule and pragma as ours, but slight enhancements in QoR are attributed to how Merlin handles memory transfer.
{adjustwidth}
-1in-1in
GF/s Perf. Improvement
Kernel TR T Sisyphus AutoDSE NLP-DSE
atax 1 D 1.96 0.99x 1.00x
atax 0 D 1.86 0.98x 1.00x
heat-3d - - 131.33 35.09x 0.99x
jacobi-2d - - 128.83 13.06x 6.60x
2mm 1 D,P 164.34 403.55x 1.33x
2mm 0 D,P 162.24 775.26x 1.36x
bert_100_64 1 D,P 88.30 1.08x 1.03x

### v4

@@ 4.3. Objective Function
We structure the objective function akin to the approach described in (Pouget et al., 2024b, a), tailoring it to meet the specific needs of our problem. Here, $Lat^{s}_{2}$ corresponds to the fine-grained unrolled level (level 2) for the statement $s$, $Lat^{s}_{1}$ denotes the pipeline tile that incorporates the fine-grained unrolled level (level 1), and recursively, $Lat_{0}$ encompasses $Lat_{1}$, enabling coarse-grained unrolling and facilitating on-chip data caching (level 0). The memory latency $Lat^{s}_{mem}$ for a statement $s$ corresponds to the time required to transfer the array from off-chip to on-chip and back for the arrays needed by a statement $s$. We assume the memory transfer is pipelined, allowing the transfer of one data element (with a maximum bitwidth of 512 bits) per cycle. Therefore, the latency at a given level is calculated as the array footprint divided by the burst size (bitwidth used for transferring the array), multiplied by the trip count of the loop that iterates over the function responsible for the transfer, based on the selected permutation of level 0. Additionally, if an array is fully transferred on-chip for a statement $s$ and reused by a subsequent statement $s^{’}$, no further transfers are needed. The burst size for an array must be a divisor of its last dimension. For instance, if we have an array $float\text{ }A[512][20]$, the maximum burst size of this array $A$ is 128 bits, since the last dimension, $20\times 32$ bits, can be divided by the largest power of 2, which is 128.
@@
$\left.\begin{cases}&Lat^{s}_{2}=IL_{par}+IL_{seq}\times\prod_{l\in\mathcal{L}^{red}}\times\log_{2}(TC_{l,2})\\ &Lat^{s}_{1}=Lat^{s}_{2}+II\times(TC_{l,1}-1)\\ &Lat^{s}_{0}=\prod_{l\in\mathcal{L}}\frac{TC_{l,0}}{loop_{l}\_UF}\times Lat^{s}_{1}\\ &L^{s}_{mem}=\sum_{l\in\mathcal{L}}\max_{a\in\mathcal{A}}(cache_{l}\_arr_{a}\times ft\_array_{a}\_loop_{l}\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\times TCs(perm_{s})/burst_{a})\\ &Lat_{s}=Lat^{s}_{0}+L^{s}_{mem}\end{cases}\right.$
Finally, the objective function (Eq. 20) is defined as the sum of the latencies of each loop body.
(20) $obj\_func=\sum_{s\in\mathcal{S}}Lat_{s}$
5. Post-Optimization and Usage
The NLP determines the trip count for each loop, array partitioning, and on-chip buffer sizes. Functions for loading/storing data between off-chip and on-chip memory are automatically generated with the maximum burst size. Further optimizations are also applied to improve QoR. Optimization for Non-Constant Trip Count  To optimize loops with non-constant trip counts, we implement a code transformation that preserves the NLP’s estimation while simplifying compilation for the HLS compiler. We achieve this by replacing loops with non-constant trip counts with the maximal trip count computed using PoCC (PoCC, [n. d.]). Next, we replace all statements iterated by this loop with calls to a function. For example, instead of using $C[i][i]+=B[i][i]$, we replace it with $C[i][j]=f(C[i][j],B[i][i])$, where the function $f$ ensures compliance with the constraints of the non-constant trip count loop. This function returns the computed value if the constraints are satisfied; otherwise, it simply returns the output value. Therefore, we employ these techniques to reduce compilation time and achieve designs with a good QoR. Grouping Data Transfers  In the NLP, we already account for parallel transfers of arrays when they are transferred at the same level in the code (e.g., within the same loop). However, we do not consider parallel transfers between arrays belonging to different loop bodies. In the 2mm example, we first handle the transfer of arrays required for the initial matrix multiplication, only transferring the arrays for the second multiplication after the first is complete. However, FPGA DRAM banks enable parallel transfers of different arrays. Therefore, we can optimize load and store operations during post-processing to enhance memory transfer overlap. Code Transformation for HLS  To simplify the understanding of the HLS compiler, particularly to ensure that the compiler can pipeline loops with the correct initiation interval (II), we simplify the reductions (regardless of whether we use tree reductions) in the fine-grained unrolled part (level 2). Therefore, under the pipeline we create a variable to accumulate the reduction, which we then add to the output. Additionally, we use the loop_flatten off pragma on all reduction loops above the pipeline. Although this adjustment could theoretically improve the QoR, in practice, flattening these loops can complicate the code and potentially increase the initiation interval (II) of the pipeline. Usage  Our framework is designed to be explainable and user-friendly, requiring the user to provide only the affine C/C++ code of the kernel to be optimized and the available resources. To fully benefit developers, our NLP formulation allows the template to be completely customized to the user’s needs. Each variable in the design space can be defined by the user if they have specific constraints. For example, in the gemm template (Listing 3), the user can set the permutations for level 0 while allowing the NLP to explore other parameters. Additionally, they can specify the on-chip size of matrix A, enabling exploration within the defined constraints.
@@ 6. Evaluation
The goal of this evaluation is to demonstrate how effectively our approach applies code transformations and inserts hardware directives while respecting user-defined resource constraints. We compared our method against four other frameworks: AutoDSE (Sohrabizadeh et al., 2021), NLP-DSE (Pouget et al., 2024b), HARP (Sohrabizadeh et al., 2023), and ScaleHLS (Ye et al., 2022). Since generating a bitstream without significant manual intervention is challenging for some frameworks, we first demonstrate our ability to achieve a high-quality design by utilizing all available resources—ensuring fair comparison with the other frameworks—without considering place and route. We generate HLS reports using the vitis-flow, which accounts for memory transfers between off-chip and on-chip, necessitating the inclusion of memory transfer steps in ScaleHLS’s code. Thus, we show that our designs are synthesizable by targeting the resources of a single SLR, which eliminates the need for manual partitioning. Due to the time required for bitstream generation, we restrict our evaluation to a subset of the kernels.
@@ 6.1. Setup
We conducted experiments using kernels from Polybench/C 4.2.1 (PolyBench, [n. d.]), a CNN kernel, and matrix multiplication tasks similar to those in the BERT transformer model. All computations used single-precision floating-point data for direct comparison with frameworks like AutoDSE, NLP-DSE, and ScaleHLS, using medium-sized Polybench/C datasets. We also compare our approach with HARP (Sohrabizadeh et al., 2023), which enhances HLS design exploration using a Graph Neural Network (GNN) model to predict HLS tool behavior. HARP explores this space for one hour and synthesizes the top 10 designs. We use the medium-sized Polybench kernels found in their training set with double-precision floating-point data types, aiming to deploy HARP under optimal conditions. As detailed in (Pouget et al., 2024b), HARP’s effectiveness is not universally applicable without the necessary fine-tuning or training, particularly when confronted with diverse problem sizes. We selected these problem sizes to highlight how our method performs in transforming code and inserting pragmas in cases where full unrolling is not feasible. We also demonstrate the effectiveness of our framework in tiling large tasks such as CNN and bert_3072_100, ensuring that only a small portion (a tile) of each array fits on-chip, in line with memory constraints. For the CNN kernel, the problem size and loop dimensions are the same as described in Section 2. For bert_n_m matrix multiplication, the dimensions are I = n, J = m, K = n (reduction). We used the AMD/Xilinx Vitis HLS compiler (Xilinx, 0232) to demonstrate our method’s effectiveness, generating reports with Vitis 2023.2. When enabling tree reduction, we selected the “funsafe math optimizations” option to allow commutative and associative reduction operators for logarithmic time reductions. For our hardware setup, we targeted the Xilinx Alveo U200 device running at 250 MHz. We use the commercial solver Gurobi 11.0.0 (Gurobi Optimization, LLC, 2024) with AMPL. The experiments were performed on an AMD EPYC 7V13 64-core processor, with 32 GB of memory and 4 cores allocated for each run.
@@ 6.2. Experimental Evaluation
To ensure a fair comparison, we modified kernels with non-constant trip counts for all benchmarks, as detailed in Section 5, except for ScaleHLS, which already handles such loops. We generated AutoDSE’s design space using the ds_generator command, applying all trip count factors for unroll and tile sizes. Since AutoDSE lacks array partitioning, its design space is slightly larger. AutoDSE was run with a 1,000-minute exploration timeout and a 180-minute HLS synthesis timeout. For NLP-DSE, we followed the original paper’s parameters. HARP was run with the authors’ settings using Vitis 2021.1 and tree reduction enabled. ScaleHLS assumes that all kernel data are already located in on-chip memory. To provide a fair comparison, we modified their code to handle off-chip to on-chip memory transfers. We optimized these transfers to 512 bits per cycle, with all arrays transferred in parallel. Each array was assigned to a different DRAM bank. However, due to the extensive manual modifications required, we did not extend this comparison to all kernels. For Sisyphus evaluation, we set the Gurobi solver timeout to 14,400 seconds (4 hours) per design. By default, we assume optimistic resource reuse between non-parallel statements. If DSP utilization exceeds limits, we rerun the NLP with pessimistic assumptions (no resource reuse) and regenerate the design. The maximum partitioning constraint, $max_{part}$, was set to 1024, per AMD/Xilinx guidelines. Each synthesis starts with a C-simulation to verify correctness. We generate Vitis reports using the vitis-flow, which counts memory transfers for all frameworks and considers the resources of the entire FPGA. For bitstream generation, we consider only one SLR to eliminate any manual intervention. We consider 60% of the resources of a single SLR to quickly identify a synthesizable design, minimizing the number of times AutoDSE needs to be run.
@@ 6.3. Comparison of the Methods
6.3.1. AutoDSE, NLP-DSE and Scale-HLS with tree reduction
Table 3 provides a comparison among AutoDSE (A-DSE), NLP-DSE (N-DSE), Scale-HLS (S-HLS) and Sisyphus (ours) with tree reduction. The T column indicates the specific transformations applied, with D representing distribution, T for data tiling, and P for permutation. The column Perf. Impr. of ours framework vs. displays the performance improvement of our framework compared to the three different frameworks. Across all evaluated kernels, we consistently achieve performance that is comparable to or better than alternatives, except for the atax, and gesummv kernels, where we observe a slowdown of less than 5%. For 3mm, symm and trmm we had to apply the constraint indicating no reuse (Eq. 17), as the constraint with optimistic reuse (Eq. 16) resulted in kernels with resource over-utilization. For Gesummv, ScaleHLS is 1.41x faster. However, 91% of our design’s execution time is spent on data transfers. ScaleHLS benefits from 512-bit data transfers, which we manually applied for testing. Our framework uses 64-bit chunks, the largest bitwidth that divides the array. With 512-bit transfers, our framework could achieve 14.50 GF/s. For other memory-bound kernels like Atax and Gesummv, the slight performance differences observed with other frameworks are largely due to how Merlin handles memory transfers.
GF/s Perf. Impr. of ours framework vs.
Kernel T Ours A-DSE N-DSE S-HLS
2mm D,P 175.49 431.17x 1.42x 5.50x
3mm D,P 140.83 80.75x 1.02x 3.86x
atax D 1.96 0.99x 1.00x 1.23x
bicg D,P 1.96 1.97x 1.97x 1.18x
doitgen D,P 54.05 1.35x 2.71x 8.45x
gemm D,P 203.39 1.84x 1.55x 5.81x
gemver P 17.62 4.58x 1.74x 9.39x
gesummv D 1.96 0.99x 1.00x 0.71x
mvt P 13.24 1.70x 1.70x 1.26x
symm D,P 240.50 8.28x 8.28x 2,704.51x
syr2k D,P 254.69 5.52x 1.86x 174.47x
syrk D,P 161.20 6.55x 2.32x 358.95x
trmm D,P 148.53 8,345.10x 8,312.42x 2,376.91x
Average 108.88 683.91x 641.46x 434.78x
Geo Mean 40.13 9.17x 3.47x 16.15x
Table 3. Throughput (GF/s) Comparison of Sisyphus, AutoDSE, NLP-DSE, and ScaleHLS with Tree Reduction Using Vitis HLS
Sisyphus averages a speedup of 683.91x, 641.46x and 434.78x over AutoDSE, NLP-DSE and Scale-HLS respectively, across the evaluated kernels. In terms of geometric mean, Sisyphus achieves a speedup of 9.17x, 3.47x and 16.15x. Preserve the original schedule when necessary  For atax, Sisyphus preserves the original schedule and inserts pragmas, effectively maintaining the existing schedule while efficiently incorporating pragmas, similar to two DSEs focused on pragma insertion. Code transformation for memory-bound kernel  For mvt and gemver, both memory-bound kernels, effective optimization entails loop permutations to prevent redundant array transfers and accommodate array partitioning within the constraints of the AMD/Xilinx compiler’s 1024 limit. Sisyphus adeptly manages these loop permutations and pragma insertions, guaranteeing optimized performance for these kernels. For bicg, fully distributing and arranging the loops appropriately can enhance parallelization capabilities. Allow to achieve a parallelism not achievable without loop transformation  For doitgen, symm, syrk, syr2k, and trmm, the original schedule imposes constraints on the achievable level of parallelism. This limitation highlights the performance enhancements made possible by Sisyphus. We observe QoR improvements for these kernels because the code transformations enable a combination of loop unrolling that cannot be achieved even with maximal loop distribution, as each loop can be partially unrolled.
6.3.2. AutoDSE and NLP-DSE without tree reduction
Transformations without tree reduction, which may be needed to preserve floating-point data precision, must be different since the optimizations differ when tree reductions are not involved. Table 4 provides a comparison among AutoDSE (A-DSE), NLP-DSE (N-DSE) and Sisyphus (ours) without tree reduction. Sisyphus achieves an average speedup of 676.03x over AutoDSE and 567.53x over NLP-DSE across the evaluated kernels without tree reduction. The geometric mean speedup is 18.50x and 4.80x, respectively. On average, the performance improvement without tree reduction is greater than with tree reduction, particularly for gemm (Listing 2) and gemver. This is because, without tree reduction, optimizing the reduction loop becomes challenging, making code transformations even more necessary. However the difference observed between syrk and syr2k with tree reduction is due to the HLS compiler achieving an II=2 instead of the expected II=1. However, without tree reduction, the compiler successfully achieves II=1.
GF/s Perf. Impr. Ours vs.
Kernel T Ours A-DSE N-DSE
2mm D,P 171.07 817.42x 1.44x
3mm D,P 140.11 511.64x 1.04x
atax D 1.86 0.98x 1.00x
bicg D,P 1.86 1.92x 1.93x
doitgen D,P 54.05 5.46x 2.71x
gemm D,P 210.61 9.25x 9.57x
gemver P 15.01 37.28x 8.73x
gesummv D,P 1.81 0.97x 0.97x
mvt P 9.75 1.42x 1.42x
symm D,P 240.55 8.28x 7.32x
syr2k D,P 429.34 17.61x 3.14x
syrk D,P 333.03 13.54x 4.78x
trmm D,P 130.90 7,362.68x 7,333.85x
Average 133.84 676.03x 567.53x
Geo Mean 41.62 18.50x 4.80x
Table 4. Throughput (GF/s) Comparison of Sisyphus, AutoDSE and NLP-DSE without Tree Reduction Using Vitis HLS
6.3.3. AutoDSE and NLP-DSE for Bert and CNN
Table 5 provides a comparison among AutoDSE (A-DSE), NLP-DSE (N-DSE) and Sisyphus (ours). The TR column denotes the status of the tree reduction option. For bert_100_64, we observed a minor decrease in QoR, attributed to the NLP selecting a configuration where the pipeline was not applied as expected. Code transformation with tiling  The selection of loop order and tile size is critical for determining which array sizes, particularly for large problem size like bert_3072_100 and CNN layers, should be cached on-chip and where they should be transferred. In these instances, the NLP aims to enhance data reuse and minimize unnecessary memory transfers while identifying the theoretical optimal balance with parallelism.
GF/s Perf. Impr. Ours vs.
Kernel TR T Ours A-DSE N-DSE
bert_100_64 1 D,P 81.22 0.99x 0.95x